OpenStudy (anonymous):

If 3x2 – x + 9 = 0, what are the values of x?

hartnn (hartnn):

do you know quadratic formula ?

OpenStudy (anonymous):

this can help you http://www.purplemath.com/modules/quadform.htm

OpenStudy (uri):

3x2 – x + 9 = 0 6-x+9=0 x=-9-6 x=-15

OpenStudy (anonymous):

just plug values for a b and c into equation

hartnn (hartnn):

Compare your quadratic equation with \(ax^2+bx+c=0\) find \[a=...?\\b=...?\\c=...?\\\] \[ \\ \sqrt{(b^2-4ac)}=...?\] then the two roots of x are: \(\huge{x_{1,2}=\frac{-b \pm \sqrt{b^2-4ac}}{2a}}\)

OpenStudy (uri):

This is a type of linear equation.

hartnn (hartnn):

linear ? \(3x^2 \implies \) quadratic

OpenStudy (uri):

That is 3x2 not 3^2,Isn't it? @imbreanna